OpenStudy (anonymous):

@dan815 Can you see if I got this right?

OpenStudy (anonymous):

4. The stadium is in need of a medical pool to attend to injured or sick whales. The pool will be located adjacent to the holding tanks. Your job is to design a medical pool using a 3-dimensional polyhedral shape you learned about in this module. For safety reasons, the tank can be no more than one-fourth the volume of a holding tank. You will need to submit a 2-dimensional drawing, the dimensions, the volume of your pool, and proof that the pool's volume is no more than one-fourth the volume of a holding tank to your instructor. For the medical pool, I decided to use a pool the shape of a cylinder to serve for the medical pool. Now we have to see if the volume of the pool is no more than one-fourth the volume of the holding tank. We will do so using this formula: V = (3.14) (r2) (h). Before we do this though, we must first find the radius of our cylinder, and we will do so by taking our diameter of 6, and dividing it by two, which gives us 3. This means that the cylinder has a radius of 3. Now that we know this, lets see if it is no more than one-fourth the volume of the holding tanks. V = (3.14) (r2) (h). V = (3.14) (32) (36). V = (3.14) (9) (36). V = (3.14) (324). V = 1017.36 Now that we found the volume of the medical tank, let's see if it is no more than one-fourth the volume of the holding tanks. In order to do so, we will tank the volume of one of the tanks and divide it by four. 42,390 / 4 = 10597.5 As you can see the volume of the medical tank is smaller by 9580.14 feet, which means that this tank will be safe and will work in this scenario.

OpenStudy (dan815):

looks good